a) How long will it take Caleb to save $5,000 for his first car if he deposits $150 at the end of every three months, and his savings account earns interest at 6¾% compounded monthly. Answer in years and months. b) During this time, how much interest would he have earned in his account? 


1
Expert's answer
2020-11-17T16:26:37-0500

a) "5,000=150*\frac{(1+\frac{0.0675}{12})^3*((1+\frac{0.0675}{12})^{3*n}-1)}{(1+\frac{0.0675}{12})^3-1}"

n-number of quarters

"n=26,28=6\ years\ and\ 7\ months"

b)"150*26=3,900" sum of all payments

"5,000-3,900=1,100" interest he have earned